OpenStudy (bbb911):

This is Geometry: Suppose ^ABC ~ ^ JKL with BC =28, JK =12, and KL = 42. Find the perimeter of ^ABC if the perimeter of ^JKL is 72.

OpenStudy (anonymous):

Bc/Kl=28/42=2/3 so preimeter ABC/peri JKL=2/3 ABc=48

OpenStudy (bbb911):

:)